# Entanglement enhancement induced by noise in inhomogeneously monitored systems

### v1

@@ I Introduction
Indeed, the competition between a coherent drive and dissipative processes in many-body quantum systems has been extensively studied in recent years. It has been showed to lead to a plethora of non-equilibrium phases and to transitions between such phases. Important examples are dissipative phase transitions [43, 53, 44, 45, 54, 46, 55, 47, 48, 49, 50, 56, 42, 57, 58, 42, 59, 51, 60, 61, 62, 63, 52, 64], measurement induced phase transitions [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91], transient dynamical phases [92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103], non-equilibrium phases [104, 105, 106, 107, 108, 109, 110, 111] and much more.
Here we present results that suggest that noise can also be exploited to enhance the effect of coherent processes. We investigate this phenomenon within a simple toy model of free fermions [112, 113, 114, 115] consisting of two coupled chains: a system chain and an ancilla chain, each subjected to its own independent noise. We model the effect of noise (or any error or decoherence process) as arising from random projective measurements acting on the system via the Kraus representation. The model is similar to the one studied in a previous work [116] coauthored by some of us. In that work only the ancilla chain was subject to random projective measurements of the occupation number. For certain values of the ancilla hopping strength, the entanglement within the system was enhanced with stronger ancilla noise.
@@
In particular, we find two regimes where the system exhibits different behaviors depending on the ancilla dynamics. In the slow dynamics (small hopping) regime, any noise on the ancilla is detrimental to entanglement in the system. In the fast dynamics (large ancilla hopping) regime, the entanglement in the system is enhanced by noise acting on the ancilla, which protects the entangled phase, even in the presence of strong noise acting on the system. Such effect is a consequence of both the monogamy of entanglement and memory effects introduced by the fast dynamics of the ancilla, which leads the system to experience an effective (correlated) non-Markovian noise [121, 122]. This is related to recent findings that non-Markovian noise [85, 116, 123], or long-range correlated noise [124, 125] can enhance the entanglement; indeed, entanglement creation via correlated dissipation was proposed and realized [126, 127] in few-body systems.
@@ II Model
FIG. 3: Colormap of the negativity in the inner chain for a half bipartition ¯$\overline{\mathcal{E}}_{\frac{L}{2}}$ for $L=32$ for selected values of $t_{2}$. The negativity is plotted as function of the measurement probability in the outer chain $p_{2}$ and the measurement probability in the inner chain $p_{1}$. The red lines indicate the contour lines where ¯$\overline{\mathcal{E}}_{\frac{L}{2}}=2.5$, while the arrows express the gradient of the negativity.
@@
The dynamics of the monitored free fermionic ladder is investigated within the quantum trajectory approach [117, 118, 119, 120]. By simulating the evolution of the ladder for a number $N_{traj}$ of different times, the stochasticity of the measurement outcomes along each trajectory gives rise to an ensemble of different quantum trajectories. For large enough times $\tau>\tau_{st}$, the negativity that evolves along a single trajectory saturates to a steady state value around which it fluctuates. For a given trajectory $\alpha$ we define the steady state average of the negativity as
@@
which is performed over the next $m$ steps after $\tau_{st}$. A safe estimate for $\tau_{st}$ for the system sizes considered in this work is $\tau_{st}=150$ for $L\leq 64$ and $\tau_{st}=250$ for $L>64$ [116]. For concreteness, in our study we have used $m=5$, $\tau_{st}=250$ and $N_{traj}=150$.
@@ III.1 Negativity and entanglement susceptibility
which again we study as function of $p_{1}$ and $p_{2}$ for the same values of $t_{2}$ as in Fig. 3, and for a system of size $L=32$. The behavior of the susceptibility is shown in Fig. 4.
@@
This effect is highly dependent on the value of the hopping in the ancilla. An intuitive explanation is due to the principle of monogamy of entanglement, which states that only a finite amount of entanglement can be shared among different partitions of a system. In our case, we have the $A$ and $B$ partitions of the system chain and the ancilla chain. When $t_{2}\gg t_{1}$, entanglement spreading within the ancilla is more efficient than entanglement spreading within the system; by virtue of the monogamy of entanglement the negativity between $A$ and $B$ is suppressed because most of entanglement is contained within the ancilla. However, measuring the ancilla with stronger $p_{2}$ decreases its correlations and makes more entanglement available to be shared between $A$ and $B$, explaining why the negativity increases with $p_{2}$. This effect does not appear for $t_{2}\approx t_{1}$, since in this regime the entanglement content within the ancilla and within the system is comparable and thus measuring the ancilla does not free up much entanglement, but rather destroys correlations within the system in an indirect way via the coupling $t_{12}$.
Another possible factor contributing to this effect is the non-Markovianity introduced by the ancilla dynamics. In fact, the ancilla acts as a filter on the $p_{2}$ noise, which is then experienced by the system as a noise with memory effects. In particular, the non-Markovianity of the system evolution has been quantified in Ref. [116] and has been showed to be larger for higher values of $t_{2}$. Large memory effects lead to a backflow of information from the environment into the system, leading to a revival of correlations [85] and can effectively act as antinoise [123]. Moreover, correlated noise in space [125, 124, 146] has been shown to substantially increase entanglement, depending on the correlation length of the noise. In our model, the noise is correlated in time due to the ancilla dynamics, with $t_{2}$ acting as a control parameter of the correlation time/length. The results found so far, then suggest that temporal correlations induced by non-Markovianity have a similar effect to spatial correlations of the noise, by increasing the entanglement content in the system.
@@ III.2 Scaling analysis
The color maps in Fig. 3 only show how much entanglement is contained in the system, but do not determine if the system entanglement is in an area law phase, a logarithmic law phase, or a volume law phase. In order to quantify this property we need to perform a study of the scaling properties of the negativity with the system size $L$. We calculate the negativity ¯$\overline{\mathcal{E}}_{L/2}$ for several system sizes $L\in(8,16,24,32,48,64,80,96,128,160,192,256)$ as function of $p_{2}$. We investigate four different regimes: regime of weak ($p_{1}=0.2$) and strong $(p_{1}=0.6)$ noise on the system; regime of slow ($t_{2}=1.0$) and fast ($t_{2}=1.0$) dynamics of the ancilla. The results are reported in Fig. 5 for $p_{2}\in(0,0.2,0.4,0.6,0.8,1)$.
When the dynamics of the ancilla is slow, and the measurement rate of the system is low ($t_{2}=1.0,p_{1}=0.2$, panel a) we observe two different behaviors. At small $p_{2}$ the entanglement is clearly logarithmic in $L$ – as expected for weakly measured non-interacting systems [114, 116, 113] – while at larger values of $p_{2}$ it saturates to a constant for large enough sizes, indicating an area law. As the measurement rate of the system $p_{1}$ becomes stronger ($t_{2}=1.0,p_{1}=0.6$, panel b), instead, the system appears to always be in an area law – the results at $p_{2}=0$ suggest a saturation at large system sizes with finite size logarithmic corrections. Such behavior of the system at small values of $t_{2}$ is reasonable and expected: there is no entanglement monogamy effect, and an increase in either $p_{1}$ or $p_{2}$ decreases the entanglement, causing the system to switch from logarithmic to area law.
@@ III.3 Transition induced by noise
FIG. 7: (a)-(b) Plot of $c_{\textrm{eff}}$ as function of $1/L$ for $t_{2}=1$ (a) and $t_{2}=5$ (b), $p_{1}=0.2$ and $p_{2}=0.1$ (blue) $p_{2}=0.7$ (red). The circles represent the values of $c_{\textrm{eff}}$ as obtained from the fitting procedure and the shaded region is the associated error. The dashed lines represent a polynomial fit of the second order $c_{\textrm{eff}}=c_{0}+c_{1}/L+c_{2}/L^{2}$. The red curve in (a) corresponds to the area law phase, correctly showing that $c_{\textrm{eff}}$ goes to zero in the thermodynamic limit $1/L\rightarrow\infty$. The other curves correspond to the logarithmic phase and exhibit a non zero value of $c_{\textrm{eff}}$ in the thermodynamic limit. (c) Finite size scaling analysis of $c_{\textrm{eff}}(p_{2},L)$ for $t_{2}=1$ and $p_{1}=0.2$. The data show a good collapse with exponents $\nu=3.5$ and $\zeta=0.0014$, and the critical probability is $p_{2c}\approx 0.26$.
We observe that $c_{\textrm{eff}}$ represents a good order parameter for the transition between logarithmic phase and area law phase (where it vanishes). Therefore, it is interestingto perform a finite size scaling analysis to check if the curves of $c_{\textrm{eff}}(p_{2},L)$ exhibit a collapse; such analysis is also useful to quantitatively determine the critical probability $p_{2c}$. We fit the behavior of $c_{\textrm{eff}}(p_{2},L)$ with the function $c_{\textrm{eff}}(p_{2},L)=L^{\zeta/\nu}f(L^{1/\nu}(p_{2}-p_{2c}))$, where $p_{2c}$ is the transition probability, while $\nu$ and $\zeta$ are the critical exponents. The results are reported in Fig. 7, panel (c), where the various curves associated to different $L$ nicely collapse onto one curve, confirming once again that $p_{2}$ tunes a transition between logarithmic and area law. We find that $p_{2c}\approx 0.26$, $\nu\approx 3.5$ and $\zeta=0.0014$.

### v2

@@ I Introduction
Indeed, the competition between a coherent drive and dissipative processes in many-body quantum systems has been extensively studied in recent years. It has been showed to lead to a plethora of non-equilibrium phases and to transitions between such phases. Important examples are dissipative phase transitions [43, 53, 44, 45, 54, 46, 55, 47, 48, 49, 50, 56, 42, 57, 58, 42, 59, 51, 60, 61, 62, 63, 52, 64], measurement induced phase transitions [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97], transient dynamical phases [98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109], non-equilibrium phases [110, 111, 112, 113, 114, 115, 116, 117] and much more.
Here we present results that suggest that noise can also be exploited to enhance the effect of coherent processes. We investigate this phenomenon within a simple toy model of free fermions [118, 119, 120, 121, 122, 123, 124, 125, 126] consisting of two coupled chains: a system chain and an ancilla chain, each subjected to its own independent noise. We model the effect of noise (or any error or decoherence process) as arising from random projective measurements acting on the system via the Kraus representation. The model is similar to the one studied in a previous work [127] coauthored by some of us. In that work only the ancilla chain was subject to random projective measurements of the occupation number. For certain values of the ancilla hopping strength, the entanglement within the system was enhanced with stronger ancilla noise.
@@
In particular, we find two regimes where the system exhibits different behaviors depending on the ancilla dynamics. In the slow dynamics (small hopping) regime, any noise on the ancilla is detrimental to entanglement in the system. In the fast dynamics (large ancilla hopping) regime, the entanglement in the system is enhanced by noise acting on the ancilla, which protects the entangled phase, even in the presence of strong noise acting on the system. Such effect is a consequence of both the monogamy of entanglement and memory effects introduced by the fast dynamics of the ancilla, which leads the system to experience an effective (correlated) non-Markovian noise [132, 133]. This is related to recent findings that non-Markovian noise [88, 124, 127, 134], or long-range correlated noise [135, 136] can enhance the entanglement; indeed, entanglement creation via correlated dissipation was proposed and realized [137, 138] in few-body systems.
@@ II Model
FIG. 3: Colormap of the negativity in the inner chain for a half bipartition ¯$\overline{\mathcal{E}}_{\frac{L}{2}}$, plotted as function of the measurement probability in the outer $p_{2}$ and inner chain $p_{1}$, for $L=32$ and for selected values of $t_{2}$. The red contour lines indicate where ¯$\overline{\mathcal{E}}_{\frac{L}{2}}=2.5$, chosen to show qualitatively the low and high entanglement regions;
the arrows express the gradient of the negativity.
@@
The dynamics of the monitored free fermionic ladder is investigated within the quantum trajectory approach [128, 129, 130, 131]. We simulate the evolution of the ladder for a number $N_{traj}$ of different times for a random |$\ket{\Psi(0)}$. The stocasticity of the measurement outcomes along each trajectory gives rise to an ensemble of different quantum trajectories. For large enough times $\tau>\tau_{st}$, the negativity that evolves along a single trajectory saturates to a steady state (independent of the initial conditions [147, 127]) value around which it fluctuates. For a given trajectory $\alpha$ we define the steady state average of the negativity as
@@
which is performed over the next $m$ steps after $\tau_{st}$  [^1] We note that fluctuations around the steady state are uncorrelated for each trajectory, so that averaging over trajectories gets rid of these fluctuations, without needing to average over the last time steps for each trajectory.. A safe estimate for $\tau_{st}$ for the system sizes considered in this work is $\tau_{st}=150$ for $L\leq 64$ and $\tau_{st}=250$ for $L>64$ [127]. For concreteness, in our study we have used $m=5$, $\tau_{st}=250$ and $N_{traj}=150$.
@@ III.1 Negativity and entanglement susceptibility
which we study as function of $p_{1}$ and $p_{2}$ for the same values of $t_{2}$ as in Fig. 3, and for a system of size $L=32$. The behavior of the susceptibility is shown in Fig. 4.
@@
This effect is highly dependent on the value of the hopping in the ancilla. An intuitive explanation is due to the principle of monogamy of entanglement, which states that only a finite amount of entanglement can be shared among different partitions of a system. In our case, we consider partitions of the ancilla and the $A$ and $B$ partitions of the system. When $t_{2}\gg t_{1}$, entanglement spreading within the ancilla is more efficient than within the system; by virtue of the monogamy of entanglement the negativity between $A$ and $B$ is suppressed because most of entanglement is shared between partitions of the ancilla; we also point out that also entanglement between the system and ancilla themselves is suppressed because $t_{2}\ll t_{12}$. However, measuring the ancilla with stronger $p_{2}$ decreases its correlations and makes more entanglement available to be shared between $A$ and $B$, explaining why the negativity increases with $p_{2}$. This effect does not appear for $t_{2}\approx t_{1}$, since in this regime the entanglement content within the ancilla and within the system is comparable and thus measuring the ancilla does not free up much entanglement, but rather destroys correlations within the system in an indirect way via the coupling $t_{12}$.
Another possible factor contributing to this effect is the non-Markovianity introduced by the ancilla dynamics. In fact, the ancilla acts as a filter on the $p_{2}$ noise, which is then experienced by the system as a noise with memory effects. In particular, the non-Markovianity of the system evolution has been quantified in Ref. [127] and has been showed to be larger for higher values of $t_{2}$. Large memory effects lead to a backflow of information from the environment into the system, leading to a revival of correlations [88] and can effectively act as antinoise [134]. Moreover, correlated noise in space [136, 135, 158] has been shown to substantially increase entanglement, depending on the correlation length of the noise. A similar enhancement was found for noise arising from certain regimes of weak spatial and temporal disorder in [89, 124]. In our model, the noise is correlated in time due to the ancilla dynamics, with $t_{2}$ acting as a control parameter of the correlation time/length. The results found so far, then suggest that temporal correlations induced by non-Markovianity have a similar effect to spatial correlations of the noise, by increasing the entanglement content in the system.
@@ III.2 Scaling analysis
The color maps in Fig. 3 only show how much entanglement is contained in the system, but do not determine if the system entanglement is in an area law phase, a logarithmic law phase, or a volume law phase. In order to quantify this property we need to perform a study of the scaling properties of the negativity with the system size $L$. We calculate the negativity ¯$\overline{\mathcal{E}}_{L/2}$ for several system sizes $L\in(8,16,24,32,48,64,80,96,128,160,192,256)$ as function of $p_{2}$. We investigate four different regimes: regime of weak ($p_{1}=0.2$) and strong $(p_{1}=0.6)$ noise on the system; regime of slow ($t_{2}=1.0$) and fast ($t_{2}=5.0$) dynamics of the ancilla. The results are reported in Fig. 5 for $p_{2}\in(0,0.2,0.4,0.6,0.8,1)$.
When the dynamics of the ancilla is slow, and the measurement rate of the system is low ($t_{2}=1.0,p_{1}=0.2$, panel a) we observe two different behaviors. At small $p_{2}$ the entanglement is clearly logarithmic in $L$, and at larger values of $p_{2}$ it saturates to a constant for large enough sizes, indicating an area law. This is expected for weakly measured non-interacting systems [120, 89, 127]; while analytic calculations [119] suggest that the logarithmic scaling is a finite size effect, our model has a different underlying unitary dynamics, and it may still display a logarithmic phase. As the measurement rate of the system $p_{1}$ grows ($t_{2}=1.0,p_{1}=0.6$, panel b), the system appears instead to always be in an area law – the results at $p_{2}=0$ suggest a saturation at large system sizes with finite size logarithmic corrections. Such behavior of the system at small values of $t_{2}$ is reasonable and expected: there is no entanglement monogamy effect, and an increase in either $p_{1}$ or $p_{2}$ decreases the entanglement, causing the system to switch from logarithmic to area law.
@@ III.3 Transition induced by noise
FIG. 7: (a)-(b) Plot of $c_{\textrm{eff}}$ as function of $1/L$ for $t_{2}=1$ (a) and $t_{2}=5$ (b), $p_{1}=0.2$ and $p_{2}=0.1$ (blue) $p_{2}=0.7$ (red). The circles represent the values of $c_{\textrm{eff}}$ as obtained from the fitting procedure and the shaded region is the associated error. The dashed lines represent a polynomial fit of the second order $c_{\textrm{eff}}=c_{0}+c_{1}/L+c_{2}/L^{2}$. The red curve in (a) corresponds to the area law phase, correctly showing that $c_{\textrm{eff}}$ goes to zero in the thermodynamic limit $1/L\rightarrow\infty$. The other curves correspond to the logarithmic phase and exhibit a non zero value of $c_{\textrm{eff}}$ in the thermodynamic limit. (c) Finite size scaling analysis of $c_{\textrm{eff}}(p_{2},L)$ for $t_{2}=1$ and $p_{1}=0.2$. The data show a good collapse with exponents $\nu=3.5\pm 1.8$ and $\zeta=0.001\pm 0.006$, and the critical probability is $p_{2c}=0.26\pm 0.09$.
We observe that $c_{\textrm{eff}}$ represents a good order parameter for the transition between logarithmic phase and area law phase (where it vanishes). Therefore, it is interesting to perform a finite size scaling analysis [163, 164] to check if the curves of $c_{\textrm{eff}}(p_{2},L)$ exhibit a collapse; such analysis is also useful to quantitatively determine the critical probability $p_{2c}$. We fit the behavior of $c_{\textrm{eff}}(p_{2},L)$ with the function $c_{\textrm{eff}}(p_{2},L)=L^{\zeta/\nu}f(L^{1/\nu}(p_{2}-p_{2c}))$, where $p_{2c}$ is the transition probability, while $\nu$ and $\zeta$ are the critical exponents. The results are reported in Fig. 7, panel (c), where the various curves associated to different $L$ nicely collapse onto one curve, confirming once again that $p_{2}$ tunes a transition between logarithmic and area law. We find that $p_{2c}=0.26\pm 0.09$, $\nu\approx 3.5\pm 1.8$ and $\zeta=0.001\pm 0.006$. Except for the critical probability, both $\nu$ and $\zeta$ have large uncertainties.